Ibarra Corporation uses the FIFO method in its process costing system. The first processing department, the Forming Department, started the month with 17,600 units in its beginning work in process inventory that were 20% complete with respect to conversion costs. The conversion cost in this beginning work in process inventory was $35,250. An additional 113,000 units were started into production during the month and 116,000 units were completed and transferred to the next processing department. There were 14,600 units in the ending work in process inventory of the Forming Department that were 30% complete with respect to conversion costs. A total of $609,123 in conversion costs were incurred in the department during the month. The cost per equivalent unit for conversion costs for the month is closest to:
Answer #1
Units started and completed = Total units competed - Units in the beginning work in process =116000-17600
98400 units

Calculation using FIFO method for conversion cost :

Beginning work in process units*(100%-%completed) =17600*(100%-20%) 14080
Started and completed*100% =98400*100% 98400
Ending work in process units*%completed =14600*30% 4380
Equivalent units 116860
Cost per equivalent unit = Total conversion cost for the month/Equivalent units of conversion cost =$609123/116860units
$5.21